OpenStudy (anonymous):

What is the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). Please help, struggling to finish FLVS!

OpenStudy (anonymous):

let the line be Y=MX+C,where M is slope and C is intercept as it passes through (6,-1) so -1=6M+C as it perpendicular to line 3x+y=7=>y=-3x+7 so as we know 2 lines perpendicular to each other has sum of slopes as zero. i.e. M-3=0 or,M=3 hence -1=6*3+C or C=-19 so the required line is Y=3X-19

OpenStudy (anonymous):

Thank you for your help, Here are the answers, I think the closest one is the last one ...? y = x − 3 y = x + 17 y = −3x − 3 y = −3x + 17

OpenStudy (anonymous):

@torilaurennn i might have done some flaws for which m not getting exact answer but I tell u a trick to reduce the no. of options. if 2 lines are parallel then their slopes will be (-1/n)th times of other if n is 1st slope. if 2 lines are perpendicular then their slopes will be same in magnitude opp. in sign

OpenStudy (anonymous):

so a and b can't be the right options

OpenStudy (anonymous):

Okay, thank you so much

OpenStudy (anonymous):

if the lines are parallel the magnitudes of the slopes are equal if the lines are perpendicular the product of the magnitudes of the two slopes equals to -1 above answer has mentioned it in a wrong way